3. Find the volume and lateral surface area of the frustum of a regular square pyramid whose altitude is 38 cm and whose base edges are 10 cm and 20 cm respectively.
1
Expert's answer
2020-12-07T20:14:39-0500

We know that ,if "a,b" and "h" are the base edge,top edge and height of a frustum of a regular square pyramid,then formula for volume and lateral surface area are -

Volume"(V) =(1\/3)(a^2+b^2+ab)\u00d7h"

Lateral surface area"(F)=2(a+b)\sqrt(((a-b)/2)^2+h^2)"

Now according to the given problem,

"a=20 cm" , "b=10cm" ,"h=38cm"

Therefore required volume of the frustum of square pyramid is

"V=(1\/3)\u00d7((20)^2+(10)^2+(20\u00d710))\u00d738"

"=(1\/3)\u00d7(400+100+200)\u00d738" "cm^3"

"=(26600\/3)" "cm^3"

and Lateral surface area is

"F=2×(20+10)×\sqrt(((20-10)/2)^2+38^2)"

"=2×30×\sqrt(25+1444)" "cm^2"

"=60×\sqrt1469" "cm^2"